OpenStudy (anonymous):

5/3 = 3x/2 full working plase

OpenStudy (anonymous):

mutltiply both sides by 2, then divide by 3. or do it in one step \[\frac{5}{3}=\frac{3x}{2}\] \[\frac{2}{3}\times \frac{5}{3}=x\] \[\frac{10}{9}=x\]